Problem 4-55 (algorithmic) Question Help * Leon and Heidi decided to invest $3,500 annualy for only the first eight years of their marriage. The first payment was made at age 20. If the annual interest rate is 9%, how much accumulated interest and principal will they have at age 70? The accumulated interest and principal will equal (Round to the nearest dollar)
Answer #1

First payment is made at age of 20 years. Eighth payment will be made at the age of 27

First, we calculate the amount accumulated at the age of 27

Annual contribution =R=$3500

Number of periods=n=8

Interest rate=i=9%

Accumulated amount at the age of 27=R*(F/A,9%,8)

(FA990,8)-1 (1 + 9%) 8-1 -11.028474 9%  

Accumulated amount at the age of 27=R*(F/A,9%,8)=3500* 11.028474=$38,599.66

Now we calculate the final accumulated amount.

This amount will earn interest for 43 (70-27) years.

Amount accumulated at the age of 70=38599.66(F/P,9%,43)=38599.66*(1+9%)^43=1,570,084.01

Principal amount=3500*8=$28,000

Accumulated Interest=1570084-28000=$1,542,084